A 0.1 significance level is used for a hypothesis test of the claim that when parents use a particular method of gender selecb. What is the value of a? a = (Type an integer or a decimal.)c. What is the sampling distribution of the sample statistic? Student (t) distribution Normal distributiond. Is the test two-tailed, left-tailed, or right-tailed? Left-tailed Right-tailed Two-tailede. What is the value of the test statistic? The test statistic is . (Type an integer or a decimal.)f. What is the P-value? The P-value is (Round to four decimal places as needed.)g. What are the critical value(s)? The critical value(s) is/are (Round to two decimal places as needed. Use a comma to separah. What is the area of the critical region? The area is (Round to two decimal places as needed.)
Answer #1

This is test for the binomial proportion. There are two outcomes either being boy or girl. Since n > 30 and np0 > 10, we can use the normal approximation.

n= 169 null proportion p0= 0.5

We want to test if the proportion for baby girl is greater than 0.5. So this is right one tailed test.

Test

a)

Ho :P 0.5

Hp > 0.5

b.

level of significance = \alpha

\alpha = 0.1

c.

The sampling distribution used will be

Normal distribution

d)

Since we are testing if greater than null proportion

Right tailed test

e)

n = 169 x: no of girls x= 91  

1596043628521_blob.png = x /n

=7/ 13 = 0.5385

Test Stat = р ро po(1 - po)/n

Where the null proportion = 0.5

Test Stat = 1

f)

p-value = P(Z > |Test stat|)

=P(Z > 1)

=1 - P( Z < 1)

=1 - 0.84135 ....................using the normal distribution tables

p-value = 0.15866

g)

C.V. = Za

=Z0.1

C.V. = 1.2816

h)

The area is test stat > C.v.

Here Test stat < C.V.

we do not reject the null hypothesis at 10%. there is insufficient evidence to conclude that the proportion os girls is greater than 0.5.